OpenStudy (anonymous):

Assembly of a television set in the factory is set at 45 minutes. Some workers take a shorter time (t) at this while some take longer. However, each varies no more than 8 minutes. Which inequality below represents this scenario? |t - 8| less than or greater to 45 |t + 8| greater than or equal to 45 |t + 45| greater than or equal to 8 |t - 45| less than or greater to 8

OpenStudy (anonymous):

t+8 gt or eq 45 at the very fastest times t=37, thus t+8=45 at slower times lets say t=40 t+8=48 ..even slower times.. t=48, t+8=56..